Exercise 15-10

For a recent 2-year period, the balance sheet of Oriole Company showed the following stockholders’ equity data at December 31 (in millions).

2017

2016

Additional paid-in capital $ 930 $ 843
Common stock 651 642
Retained earnings 7,210 5,220
Treasury stock 1,850 945
   Total stockholders’ equity $6,941 $5,760
Common stock shares issued 217 214
Common stock shares authorized 500 500
Treasury stock shares 37 27

(a) Answer the following questions.

(1) What is the par value of the common stock? (Round par value to 2 decimal places, e.g. $3.15.)
Par value of common stock $


(2) What is the cost per share of treasury stock at December 31, 2017, and at December 31, 2016?

December 31, 2017

December 31, 2016

Cost per share of Treasury stock $

$


(b) Prepare the stockholders’ equity section at December 31, 2017. (Enter account name only and do not provide descriptive information.)

Oriole Company
Balance Sheet (in millions of dollars) (Partial)
Answer #1

Par value of common stock =Common stock /number of shares issued

   For 2017 :   651/217 =$ 3 per share

For 2016 :642/214 = $ 3 per share

so Par value = $ 3 per share

2)

December 31, 2017 December 31, 2016
Cost per share of Treasury stock   [Cost of treasury stock /number of shares in treasury] 1850/37 =$ 50 per share 945/27 = $ 35per share

b)

Oriole Company
Balance Sheet (in millions of dollars) (Partial)

Stockholders equity
Common stock 500 shares authorized ,$ 3 par,217 shares issued 651
Additional paid-in capital 930
Total paid in capital 1581
Retained earning 7210
Total paid in capital and retained earning 8791
less:Treasury stock (37 shares ) (1850)
Total stockholders equity 6941